# Slant sums of quiver gauge theories

### v1

@@ 5.4 Mirror symmetry and some conjectures
One should be able to prove this conjecture by using microlocalization of $\mathcal{A}_{x}$ to a sheaf of algebras on $\mathcal{M}_{C}$. Since $\mathcal{M}_{C}$ is not conical, one should be careful. We will return to this question in the second version of the paper.
@@
Finally, let’s note that the quantized counterpart of Conjecture 5.3 should claim that there exists a homomorphism between the corresponding quantized Coulomb branches. Then, taking modules $L$ as above, one would obtain a collection of modules over the quantization of $\mathcal{M}_{C}$ labeled by nonsingular fixed points of $\widetilde{\mathcal{M}}_{C,\nu}$. Let’s decompose $\lambda=\lambda_{1}+\ldots+\lambda_{N}$ into the sum of fundamental weights. Then, we obtain the decomposition $\mu=\mu_{1}+\ldots+\mu_{N}$ ($\mu_{i}=w(\lambda_{i})$). The module $L$ should be isomorphic to a tensor product of modules $L_{i}$ corresponding to $\lambda_{i},\mu_{i}$. For $Q$ of type ADE the corresponding modules $L_{i}$ are so-called prefundamental modules over $Y_{\mu_{i}}$.
6 Coulomb branches for ADE quivers
The goal of this section is to prove Conjectures 1.5, 5.2, and 5.3 for ADE quiver theories. Conjectures 5.2 and 5.3 can be deduced from the existing literature and our argument for the proof of Conjecture 1.5 is based on the approach used in [25]. From now on, we assume that $Q$ is of type ADE.
@@ 6.1 Realization via slices in affine Grassmannians
In this section, we recall all of the objects we need to prove conjectures above. Let $\mathfrak{g}=\mathfrak{g}_{Q}$ be the simple Lie algebra corresponding to $Q$. Let $G=G_{Q}$ be the adjoint Lie group with Lie algebra $\mathfrak{g}$ and choose a Borel subalgebra $B$ with opposite Borel $B_{-}$. Let $\mathsf{A}_{Q}$ be the maximal torus of $G$ (i.e. the intersection of $B$ and $B_{-}$) and let $\mathfrak{a}_{Q}$ be its Lie algebra. For $i\in Q_{0}$ we will denote by $\omega_{i}^{\vee},\alpha_{i}^{\vee}$ the corresponding fundamental coweight and simple coroot for $\mathfrak{g}$. Given $\mathsf{v},\mathsf{w}\in\mathbb{N}^{Q_{0}}$, set:
$\lambda=\sum_{i}\mathsf{w}_{i}\omega_{i}^{\vee},~\mu=\lambda-\sum_{i}\mathsf{v}_{i}\alpha_{i}^{\vee}.$
Following [6], we define the generalized transversal slice in the affine Grassmannian $\overline{\mathcal{W}}_{\mu}^{\lambda}$. It is the moduli space of the data $(\mathcal{P},\sigma,\phi)$, where:
@@
It follows from [6, Section 2(xi)] that $\overline{\mathcal{W}}^{\lambda}_{\mu}$ has the following matrix description:
@@
where the notation is as in [6]. The open subvariety $\mathcal{W}^{\lambda}_{\mu}\subset\overline{\mathcal{W}}^{\lambda}_{\mu}$ consists of points $g\in\overline{\mathcal{W}}^{\lambda}_{\mu}$ that lie in $G[z]z^{\lambda}G[z]$.
It follows from [6, Theorem 3.10] that we have an isomorphism:
@@
For a coweight $\eta\in\mathfrak{a}_{Q}$ of $\mathfrak{g}$ we will denote by $\eta^{*}$ the coweight $w_{0}(\eta)$ (here $w_{0}$ is the longest element of the Weyl group $W$ acting naturally on $\mathfrak{a}_{Q}$). Strictly speaking, in [6, Theorem 3.10] the author construct an isomorphism $\mathcal{M}_{C}(\mathsf{v},\mathsf{w})\simeq\overline{\mathcal{W}}_{\mu^{*}}^{\lambda^{*}}$. Note now that there exists an identification:
@@
Clearly, $\lambda=\lambda_{1}+\ldots+\lambda_{N}$. Set $\mathsf{T}_{Q}:={\mathsf{A}}_{Q}\times\mathbb{C}^{\times}_{\hbar}$. It follows from [7, Section 5] that we have a $\mathsf{T}_{Q}$-equivariant isomorphism of varieties:
@@
See [24, Lemma 2.8]. ∎
More generally, it follows from [25, Proposition 5.9] that the $\theta$-fixed points on $\widetilde{\mathcal{W}}^{\underline{\lambda}}_{\mu}$ are isolated and are in bijection with $N$-tuples of coweights $\underline{\mu}=(\mu_{1},\ldots,\mu_{N})$ such that $\mu=\mu_{1}+\ldots+\mu_{N}$ and $\mu_{i}$ is a weight of $V(\lambda_{i})$. We will denote the point corresponding to $\underline{\mu}$ by $z^{\underline{\mu}}$. In [6, 2(vi)], the authors defined the so-called multiplication morphism: ${\bf{m}}^{\underline{\lambda}}_{\underline{\mu}}\colon\overline{\mathcal{W}}^{\lambda_{1}}_{\mu_{1}}\times\ldots\times\overline{\mathcal{W}}^{\lambda_{N}}_{\mu_{N}}\rightarrow\overline{\mathcal{W}}^{\lambda}_{\mu}$ that lifts to a morphism:
@@
The morphism $\tilde{\bf{m}}^{\underline{\lambda}}_{\underline{\mu}}$ is known to be an open embedding (see [25, Proposition 5.7]). It is also known to be ${\mathsf{A}}_{Q}$-equivariant, and moreover, it becomes $\mathbb{C}^{\times}_{\hbar}$-equivariant after appropriately twisting the $\mathbb{C}^{\times}_{\hbar}$-action on the source of (19) (see [25, Section 5.4] for the details). It sends the unique $\mathsf{A}_{Q}$-fixed point of the domain to the point $z^{\underline{\mu}}\in\widetilde{\mathcal{W}}^{\underline{\lambda}}_{\mu}$.
@@
6.2 Proofs for ADE quivers

### v2

@@ 6.2 Completions and a formal version of $L$
Note that $\mathbb{W}_{\hbar}(T^{*}_{p}\mathcal{M}_{C})$ contains the “ordinary” Weyl algebra as a subalgebra. By “ordinary” we mean the one given by $\mathbb{C}[T^{*}_{p}\mathcal{M}_{C},\hbar]$ subject the relations (17). Algebra $\mathbb{W}_{\hbar}(T^{*}_{p}\mathcal{M}_{C})$ is larger than the ordinary Weyl algebra.
The action $\theta(\mathbb{C}^{\times})\curvearrowright T^{*}_{p}\mathcal{M}_{C}$ induces the decomposition
$T_{p}^{*}\mathcal{M}_{C}=T^{*,+}_{p}\mathcal{M}_{C}\oplus T^{*,-}_{p}\mathcal{M}_{C}$
into positive and negative parts. Symplectic form on $\mathcal{M}_{C}^{\mathrm{reg}}$ induces the nondegenerate pairing:
$(\,,\,)\colon T^{*,+}_{p}\mathcal{M}_{C}\times T^{*,-}_{p}\mathcal{M}_{C}\rightarrow\mathbb{C}.$
For $y\in T^{*,+}_{p}\mathcal{M}_{C}$ let $v_{y}:=(y,-)$ be the corresponding element of $T^{-}_{p}\mathcal{M}_{C}$. Let’s now define an action of $\mathbb{W}_{\hbar}(T^{*}_{p}\mathcal{M}_{C})$ on $\mathbb{C}[T_{p}^{-}\mathcal{M}_{C}][[\hbar]]$. We define the action of $x\in T_{p}^{*,-}\mathcal{M}_{C}$ and $y\in T_{p}^{*,+}\mathcal{M}_{C}$ on $f\in\mathbb{C}[T^{-}_{p}\mathcal{M}_{C}][[\hbar]]=S^{\bullet}(T^{*,-}_{p}\mathcal{M}_{C})[[\hbar]]$ by the following formulas:
$x\cdot f=xf,~y\cdot f=\hbar\partial_{v_{y}}(f).$ (18)
Lemma 6.4.
The action (18) extends to the action of $\mathbb{W}_{\hbar}(T_{p}^{*}\mathcal{M}_{C})$ on $\mathbb{C}[T_{p}^{-}\mathcal{M}_{C}][[\hbar]]$.
Proof.
It is enough to check that for every integer $i\in\mathbb{Z}$ and $a\in\mathbb{W}_{\hbar}(T^{*}_{p}\mathcal{M}_{C})_{i}$, the action of $a$ is well-defined. Let’s pick a basis $x_{1},\ldots,x_{d}$ of $T_{p}^{*,-}\mathcal{M}_{C}$ consisting of $\theta(\mathbb{C}^{\times})$-eigenvectors and let $y_{1},\ldots,y_{d}\in T^{*,+}_{p}\mathcal{M}_{C}$ be the dual basis. For $j=1,\ldots,d$ let $k_{j}$ be the $\theta$-weight of $x_{j}$. Note that $k_{j}<0$ for any $j$ as above and the $\theta$-weight of $y_{j}$ is equal to $k_{j}$.
We can write $a$ as a (possibly infinite) sum of elements
$x_{1}^{l_{1}}\ldots x_{d}^{l_{d}}y_{1}^{m_{1}}\ldots y_{d}^{m_{d}},~l_{j},m_{j}\in\mathbb{Z}_{\geqslant 0},~\sum_{j=1}^{d}l_{j}k_{j}-\sum_{j=1}^{d}m_{j}k_{j}=i$
with some coefficients from $\mathbb{C}[[\hbar]]$. In particular, for any fixed $r\in\mathbb{Z}_{\geqslant 0}$, we have only finitely many possible $l_{j},m_{j}$ as above such that $\sum_{j=1}^{d}m_{j}(-k_{j})\leqslant r$. It then follows that for any fixed vector $f\in\mathbb{C}[T_{p}^{-}\mathcal{M}_{C}]$ only finite number of elements in the sum representing element $a$ would act on $f$ by nonzero operators. So, the action of $a$ on $f$ is well-defined. ∎
Lemma 6.5.
After the identification $\mathbb{W}_{\hbar}(T^{*}_{p}\mathcal{M}_{C})\simeq\widehat{\mathcal{A}}_{\hbar,\widehat{p}}^{\mathrm{fin}}$ module $\mathbb{C}[T^{-}_{p}\mathcal{M}_{C}][[\hbar]]$ becomes isomorphic to the induced module from $\mathsf{C}_{\theta}(\widehat{\mathcal{A}}_{\hbar,\widehat{p}}^{\mathrm{fin}})=\mathbb{C}[[\hbar]]$. We will denote this induced module by $M_{\hbar}$.
Proof.
Follows from the definitions. ∎
6.3 Graded traces and their formal versions
We will now need a notion of the $D$-module of graded traces. This notion was introduced in [undefad] and further studied in [undefq]. Let us recall the construction. Let $\Sigma\subset\mathsf{a}_{Q}^{*}$ be the subset of $\mathsf{A}_{Q}$-weights that appear in $T^{*}_{p}\mathcal{M}_{C}$. Set
$\Sigma_{-}:=\{\alpha\in\Sigma\,|\,\langle\alpha,\theta\rangle<0\}.$
If $V$ is a vector space then we will denote by $V[[z]]$ the space of formal linear combinations $\sum_{d\in\operatorname{Span}_{\mathbb{Z}_{\geqslant 0}}(\Sigma_{-})}z^{d}v_{d}$, $v_{d}\in V$. The space $\operatorname{GrTr}(\mathcal{A})$ is defined as follows:
$\operatorname{GrTr}(\mathcal{A})=\mathcal{A}_{0}[[z]]/\mathbb{C}[[z]]\cdot\{ab-z^{\eta}ba\},$
where $a\in\mathcal{A}_{\eta}$, $b\in\mathcal{A}_{-\eta}$ s.t. $\eta\in\operatorname{Span}_{\mathbb{Z}_{\geqslant 0}}(\Sigma_{-})$.
The following general properties of $\operatorname{GrTr}(\mathcal{A})$ are known by [undefq].
  • Module $\operatorname{GrTr}(\mathcal{A})$ is finitely generated over $\mathbb{C}[[z]]$.
  • The natural map $H[[z]]\rightarrow\operatorname{GrTr}(\mathcal{A})$ is surjective.
  • The $z=0$ specialization of $\operatorname{GrTr}(\mathcal{A})$ is equal to $\mathsf{C}_{\theta}(\mathcal{A})$.
Lemma 6.6.
Map $\operatorname{tr}_{M_{\hbar}}(-)\colon\operatorname{GrTr}(\widehat{\mathcal{A}}_{\hbar,\widehat{p}}^{\mathrm{fin}})\rightarrow\mathbb{C}[[z]][[\hbar]]$ induces the map $\operatorname{GrTr}(\mathcal{A}_{\hbar})\rightarrow\mathbb{C}[[z]][\hbar]$ that, in turn, induces the map
$\operatorname{GrTr}(\mathcal{A})\rightarrow\mathbb{C}[[z]]$
to be denoted $\operatorname{tr}$ or $\operatorname{tr}(-)$.
Proof.
Set $\mathcal{A}_{\hbar}:=\operatorname{R}_{\hbar}(\mathcal{A})$. The map $\mathcal{A}_{\hbar}\rightarrow\widehat{\mathcal{A}}_{\hbar,\widehat{p}}$ induces the $\mathbb{C}^{\times}_{\hbar}$-equivariant map:
$\operatorname{GrTr}(\mathcal{A}_{\hbar})\rightarrow\operatorname{GrTr}(\widehat{\mathcal{A}}_{\hbar,\widehat{p}}^{\mathrm{fin}})=\operatorname{GrTr}(\mathbb{W}_{\hbar}(T^{*}_{p}\mathcal{M}_{C}))=\mathbb{C}[[z]][[\hbar]].$
Passing to $\mathbb{C}^{\times}_{\hbar}$-finite vectors we obtain the $\mathbb{C}[[z]][\hbar]$-equivariant map
$\operatorname{GrTr}(\mathcal{A}_{\hbar})\rightarrow\mathbb{C}[[z]][\hbar].$ (19)
Note now that the functional $\operatorname{tr}_{M_{\hbar}}(-)\colon\operatorname{GrTr}(\widehat{\mathcal{A}}_{\hbar,\widehat{p}}^{\mathrm{fin}})=\mathbb{C}[[z]][[\hbar]]\rightarrow\mathbb{C}[[z]][[\hbar]]$ induces the map $\mathbb{C}[[z]][\hbar]\rightarrow\mathbb{C}[[z]][\hbar]$ (use that $\operatorname{tr}_{M_{\hbar}}(1)$ lives in $\mathbb{C}[[z]]$) to be denoted by the same symbol. Composing it with (19) we obtain the desired $\mathbb{C}[[z]][\hbar]$-linear map $\operatorname{GrTr}(\mathcal{A}_{\hbar})\rightarrow\mathbb{C}[[z]][\hbar]$ that restricts to the map $\operatorname{tr}\colon\operatorname{GrTr}(\mathcal{A})\rightarrow\mathbb{C}[[z]]$ sending $1$ to $\operatorname{tr}_{M_{\hbar}}(1)$.
Lemma 6.7.
The natural map $\mathbb{C}[[z]]\rightarrow\operatorname{GrTr}(\mathcal{A})$ is an isomorphism. In particular, $\mathsf{C}_{\theta}(\mathcal{A})=\mathbb{C}$ (as algebras) and $\mathcal{O}_{\theta}(\mathcal{A})$ contains uinque irreducible object $L$.
Proof.
First of all note that the map $\mathbb{C}[[z]]\rightarrow\operatorname{GrTr}(\mathcal{A})$ is surjective, this follows from Nakayama lemma combined with the fact that $\operatorname{dim}\mathsf{C}_{\theta}(\mathcal{A})\leqslant\operatorname{dim}\mathbb{C}[\mathcal{M}_{C}^{\theta}]=1$. To see that it is injective it’s enough to note that the composition $\mathbb{C}[[z]]\rightarrow\operatorname{GrTr}(\mathcal{A})\xrightarrow{\operatorname{tr}}\mathbb{C}[[z]]$ is given by the multiplication by an invertible function hence is an isomorphism. ∎
6.4 Normalized character of $L$
Set $D:=S^{\bullet}(\mathsf{a})[[z]]$ and define the algebra structure on $D$ by the following formula (see [undefad, Section 3.2]):
$z^{\eta}x\cdot z^{\xi}y=z^{\eta+\xi}xy+\langle\xi,x\rangle z^{\eta+\xi}y=z^{\eta+\xi}(x+\langle\xi,x\rangle)y$ (20)
for all $x,y\in\mathsf{a}$, $\eta,\xi\in\operatorname{Span}_{\mathbb{Z}_{\geqslant 0}}(\Sigma_{-})$. This is a module over a (local) ring $\mathbb{C}[[z]]$.
Let $Z$ be the following vector space. It consists of formal linear combinations $\sum_{\eta\in\mathsf{a}^{*}}z^{\eta}a_{\eta}$ such that:
  • there exists a finite set $\{\kappa_{1},\ldots,\kappa_{r}\}\subset\mathsf{a}^{*}$ such that for any $\eta\in\mathsf{a}^{*}$ such that $a_{\eta}\neq 0$ we have either $\eta=\kappa_{i}$ or $\langle\kappa_{i}-\eta,\theta\rangle\in\mathbb{Z}_{>0}$ for some $i\in\{1,\ldots,r\}$,
  • for any $i\in\{1,\ldots,r\}$ and $m\in\mathbb{Z}_{>0}$ the set of $\eta\in\mathsf{a}^{*}$ such that $\langle\kappa_{i}-\eta,\theta\rangle=m$ is finite.
Algebra $D$ acts on $Z$ via the following formulas:
$x\cdot z^{\eta}=\langle\eta,x\rangle z^{\eta},~z^{\xi}\cdot z^{\eta}=z^{\xi+\eta}.$
Let $\kappa\in\mathsf{a}_{Q}^{*}$ be the composition $\mathsf{a}_{Q}\subset\mathcal{A}_{0}\twoheadrightarrow\mathsf{C}_{\theta}(\mathcal{A})=\mathbb{C}$. Space $\mathbb{C}[[z]]$ admits an action of $D$ by the formulas (note the $\kappa$-shift):
$x\cdot z^{\eta}=\langle\eta+\kappa,x\rangle z^{\eta},~z^{\xi}\cdot z^{\eta}=z^{\xi+\eta}.$ (21)
Because of the $\kappa$-shift in the definition of the $D$-module, we will denote $\mathbb{C}[[z]]$ with this $D$-module structure by $\mathbb{C}[[z]]_{\kappa}$.
Lemma 6.8.
[…]
Proof.
It follows from Lemma 6.7 that $\operatorname{tr}$ is an isomorphism. The fact that $\operatorname{tr}$ is a homomophism of $D$-modules follows from the definitions. ∎
Lemma 6.9.
The space of $D$-module homomorphisms $\operatorname{GrTr}(\mathcal{A})\rightarrow Z$ is equal to $\mathbb{C}$.
Proof.
By Lemma 6.8, we can identify $\operatorname{GrTr}(\mathcal{A})\simeq\mathbb{C}[[z]]$ so we are dealing with $D$-equivariant maps $\mathbb{C}[[z]]\rightarrow\mathbb{C}[[z]]$. Any such map is uniquely determined by the image of $1$ (use that it is $\mathbb{C}[[z]]$-linear). Let $f(z)\in\mathbb{C}[[z]]$ be the image of $1$. It must be annihilated by the action of any $x\in\mathsf{a}_{Q}$ implying that $f(z)$ is a constant as desired. ∎
Lemma 6.10.
The normalized $\mathsf{a}_{Q}$-character of $L$ is equal to the $\mathsf{A}_{Q}$-character of $\mathbb{C}[T_{p}^{-}\mathcal{M}_{C}]$.
Proof.
Let $\operatorname{tr}_{L}(-)\colon\operatorname{GrTr}(\mathcal{A})\rightarrow Z$ be the graded trace of $L$. Consider also $\operatorname{tr}(-)$ constructed in Lemma 6.6. Both of them define homomorphisms of $D$ -modules $\operatorname{GrTr}(\mathcal{A})\rightarrow Z$ so by Lemma 6.9 $\operatorname{tr}_{L}(-)=\operatorname{tr}(-)$.
Now, $\operatorname{tr}(1)=\operatorname{tr}_{M_{\hbar}}(1)$ is equal to $\operatorname{tr}_{\mathbb{C}[T^{-}_{p}\mathcal{M}_{C}][[\hbar]]}(1)$. The claim then follows from the fact that the commutator action of $\mathsf{a}_{Q}$ on $\mathcal{A}$ integrates to the standard action $\mathsf{A}_{Q}\curvearrowright\mathcal{A}$. ∎
Combining all of the above we obtain the following proposition. Let $\widetilde{\operatorname{ch}}(L)$ be the $\mathsf{a}_{Q}$-character of $L$ normalized in such a way that it starts with $1$.
Proposition 6.11.
Assuming Conjecture 1.7 holds and for $\theta>0$ we have:
$\widetilde{\operatorname{ch}}(L)=\prod_{\alpha\in\Phi^{-,\mathrm{re}}_{\mu}}\frac{1}{(1-e^{\alpha})^{\langle\alpha,\mu\rangle}},$ (22)
where $L$ is the unique irreducible object in the category $\mathcal{O}_{\theta}(\mathcal{A})$.
@@ 6.4 Normalized character of $L$
By taking into account the $\mathbb{C}^{\times}_{\hbar}$-action on $T_{p}^{-}\mathcal{M}_{C}$ we obtain a natural deformation of the normalized $\mathsf{a}_{Q}$-character of $L$ given by $\prod_{\alpha\in\Phi^{-,\mathrm{re}}_{\mu}}\prod_{s=1}^{\langle\alpha,\mu\rangle}\frac{1}{1-e^{s\hbar}e^{\alpha}}$. This is compatible with the additional “vv”-parameter in [undefaq, Theorem 1.3].
Remark 6.13.
For $Q$ being a finite type Dynkin quiver, we prove Conjecture 1.7 in Section 7.2 (by combining the argument similar to the one in [undefaf, Lemma 4.4] with [undefah, Theorem 3.1(1)]). In particular, we obtain a character formula (22) for the irreducible module over the truncated shifted Yangian $Y^{\lambda}_{w\lambda}$ for arbitrary dominant $\lambda$ and $w\in W$. Another proof of this result will appear in the upcoming paper [undefab] by Artem Kalmykov, Joel Kamnitzer, Alexis Leroux-Lapierre, and Théo Pinet. A closely related result was also obtained by Negut in [undefaq, Theorem 1.3] by completely different methods.
Remark 6.14.
Let’s note that the quantized counterpart of Conjecture 5.3 should claim that there exists a homomorphism between the corresponding quantized Coulomb branches. Then, taking modules $L$ as above, one would obtain a collection of modules over the quantization of $\mathcal{M}_{C}$ labeled by nonsingular fixed points of $\widetilde{\mathcal{M}}_{C,\nu}$. Let’s decompose $\lambda=\lambda_{1}+\ldots+\lambda_{N}$ into the sum of fundamental weights. Then, we obtain the decomposition $\mu=\mu_{1}+\ldots+\mu_{N}$ ($\mu_{i}=w(\lambda_{i})$). The module $L$ should be isomorphic to a tensor product of modules $L_{i}$ corresponding to $\lambda_{i},\mu_{i}$. For $Q$ of finite type the corresponding modules $L_{i}$ are so-called chamber modules over $Y_{\mu_{i}}$ (we follow the terminology of the upcoming paper [undefab] by Kamnitzer, Leroux-Lapierre, Pinet, and Weekes). For $\mu=\lambda,w_{0}\lambda$, the corresponding modules are so-called positive and negative prefundamental modules over $Y_{\mu_{i}}$ (as in [undefx, Section 3]).
7 Coulomb branches for finite type Dynkin quivers
The goal of this section is to prove Conjectures 1.7, 5.1, and 5.3 for finite type Dynkin quivers. Conjecture 5.1 follows by combining the results of [undefal], [undefh] and Conjecture 5.3 follows from the results of [undefai]. Our argument for the proof of Conjecture 1.7 is obtained by combining the argument similar to the one in [undefaf, Lemma 4.4] with [undefah, Theorem 3.1(1)]. From now on, we assume that $Q$ is of finite type.
@@ 7.1 Realization via slices in affine Grassmannians
In this section, we recall all of the objects we need to prove conjectures above. Recall that $\mathfrak{g}=\mathfrak{g}_{Q}$ is the simple Lie algebra corresponding to $Q$. Let $G=G_{Q}$ be the adjoint Lie group with Lie algebra $\mathfrak{g}$ and choose a Borel subalgebra $B$ with opposite Borel $B_{-}$. Let $\mathsf{A}_{Q}$ be the maximal torus of $G$ (i.e. the intersection of $B$ and $B_{-}$) and let $\mathfrak{a}_{Q}$ be its Lie algebra. Recall that for $i\in Q_{0}$ we will denote by $\omega_{i}^{\vee},\alpha_{i}^{\vee}$ the corresponding fundamental coweight and simple coroot for $\mathfrak{g}$. Given $\mathsf{v},\mathsf{w}\in\mathbb{N}^{Q_{0}}$, we define $\lambda,\mu$ as in (13). Recall that $W$ is the Weyl group of $G$ and let $w_{0}\in W$ be its longest element.
Following [undefh], we define the generalized transversal slice in the affine Grassmannian $\overline{\mathcal{W}}_{\mu}^{\lambda}$. It is the moduli space of the data $(\mathcal{P},\sigma,\phi)$, where:
@@
It follows from [undefh, Section 2(xi)] that $\overline{\mathcal{W}}^{\lambda}_{\mu}$ has the following matrix description:
@@
where the notation is as in [undefh]. The open subvariety $\mathcal{W}^{\lambda}_{\mu}\subset\overline{\mathcal{W}}^{\lambda}_{\mu}$ consists of points $g\in\overline{\mathcal{W}}^{\lambda}_{\mu}$ that lie in $G[z]z^{\lambda}G[z]$.
It follows from [undefh, Theorem 3.10] for ADE types and [undefap, Theorem 4.1] for BCFG types that we have an isomorphism:
@@
For a coweight $\eta\in\mathfrak{a}_{Q}$ of $\mathfrak{g}$ we will denote by $\eta^{*}$ the coweight $w_{0}(\eta)$ (here $w_{0}$ is the longest element of the Weyl group $W$ acting naturally on $\mathfrak{a}_{Q}$). Strictly speaking, in [undefh, Theorem 3.10], authors construct an isomorphism $\mathcal{M}_{C}(\mathsf{v},\mathsf{w})\simeq\overline{\mathcal{W}}_{\mu^{*}}^{\lambda^{*}}$. Note now that there exists an identification:
@@
Clearly, $\lambda=\lambda_{1}+\ldots+\lambda_{N}$. Set $\mathsf{T}_{Q}:={\mathsf{A}}_{Q}\times\mathbb{C}^{\times}_{\hbar}$. It follows from [undefi, Section 5] that we have a $\mathsf{T}_{Q}$-equivariant isomorphism of varieties:
@@
See [undefah, Lemma 2.8]. ∎
More generally, it follows from [undefai, Proposition 5.9] that the $\theta$-fixed points on $\widetilde{\mathcal{W}}^{\underline{\lambda}}_{\mu}$ are isolated and are in bijection with $N$-tuples of coweights $\underline{\mu}=(\mu_{1},\ldots,\mu_{N})$ such that $\mu=\mu_{1}+\ldots+\mu_{N}$ and $\mu_{i}$ is a weight of $V(\lambda_{i})$. We will denote the point corresponding to $\underline{\mu}$ by $z^{\underline{\mu}}$. In [undefh, 2(vi)], the authors defined the so-called multiplication morphism: ${\bf{m}}^{\underline{\lambda}}_{\underline{\mu}}\colon\overline{\mathcal{W}}^{\lambda_{1}}_{\mu_{1}}\times\ldots\times\overline{\mathcal{W}}^{\lambda_{N}}_{\mu_{N}}\rightarrow\overline{\mathcal{W}}^{\lambda}_{\mu}$ that lifts to a morphism:
@@
The morphism $\tilde{\bf{m}}^{\underline{\lambda}}_{\underline{\mu}}$ is known to be an open embedding (see [undefai, Proposition 5.7]). It is also known to be ${\mathsf{A}}_{Q}$-equivariant, and moreover, it becomes $\mathbb{C}^{\times}_{\hbar}$-equivariant after appropriately twisting the $\mathbb{C}^{\times}_{\hbar}$-action on the source of (27) (see [undefai, Section 5.4] for the details). It sends the unique $\mathsf{A}_{Q}$-fixed point of the domain to the point $z^{\underline{\mu}}\in\widetilde{\mathcal{W}}^{\underline{\lambda}}_{\mu}$.
@@
7.2 Proofs for finite type Dynkin quivers